OpenStudy (anonymous):

Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2.

OpenStudy (anonymous):

I don't get is -2 is slope?

OpenStudy (anonymous):

i dont understand it either :/...

OpenStudy (anonymous):

if x=-2 is linear than (2,1) can intercept ,can't not perpendicular ,

OpenStudy (anonymous):

It can be done. X = -2 is simply suggesting the line is going up and down. The perpendicular slope would be a horizontal line, or a slope of 0. y-y1 = m(x-x1) y-y1 = 0(x-x1) Solve from there.

OpenStudy (id21):

(2,1) means that x=2, y=1 If you need perpendicular line to x=-2, then you know that other line must be in form y=constant so you can say that y=1 will be that line what you are looking for, because it is line which crosses the x=-2 at point (2,1) I attach a picture to make it clear to you.

OpenStudy (anonymous):

Oh right. Didn't think of id21's easier method. Since you know the slope is zero, that means you're going to solve y = ? The "?" can be gotten straight from the y value of 2,1, which is 1. Clever thinking, bro!